SHift summary for Eve 15 July 1999

Leo Bellantoni (bellanto@fnpxsr)
Fri, 16 Jul 1999 00:02:00 CDT

Shift Summary for Eve shift, 15/Jul/99

On Shift:
Leo Bellantoni, Collin WOlfe

Hours of beam delivered at all intensities: 7.6
Hours of beam delivered at nominal intensity: 1.0 (see below)
Hours of E832 data (special runs not included): 6.8
Hours of special runs (laser scans, muon runs...): 0.2

Accounting of lost time:
Run changes
Beam down
KFREND/ repeater cyles for front end problems

========================================================================
Runs ended this shift:
Run 13915 - 24.2M events - BEAM_832
Run 13916 - 100k events - BEAM_MUON

Tapes brought to FCC this shift: QK134-139; QKB046-7. + Onsplits

========================================================================
Accesses during shift:

None

========================================================================
Overview of shift:
E832 datataking. Beam conditions mostly good but fluctuating.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13915 15:30 No beam at start of shift.
193 15:54 Beam is back, but SEM is 4e12. Throughout
the run, SEM readings vary from this low end
up to 7.2e12. One hour figure at nominal
intensity does not include this run, so is an
underestimate of number of spills at 6.8e12.
21:08 Trigger went into ERR4:T/O on cal trigger;
No permit back from stream 5. KFRENDing and
cycling repeater power fixed it, but it cost
11 spills.
13915 479 22:19 Stop run, swap tapes, start muon run. Send
following tapes to the vault:
QK134-139; QKB046-7
QKCC06, QKNL11 QKN011-012 QKA003 QKMU09 QKSM06.

13916 1 22:41 Muon run starts; however no beam. Power supply
in booster is being swapped; it is believed
that this will fix the intensity fluctuations.

13917 1 23:04 Back to hadron running

========================================================================
Follow up on Notes from Previous Shift

========================================================================
Items to Follow up on Next Shift

Banana monitoring seems to have returned to older erroneous form with
large offsets.

Keep an eye on Csi 3114, 961; Rc timing plot within bounds but not
by much.

Leo
-------------------------------------------------------------------------------
Leo Bellantoni WH13SX, FermiLab Batavia, IL 60510
Phone:(630)840-8772 Pager:(630)266-3803
"The more I use computers, the more I like pencils."
-------------------------------------------------------------------------------